OpenStudy (mlgpro):

six hundred quarts of water are drained from a swimming pool. How many pints of water were drained from the pool? one hundred fifty pints three hundred pints one thousand two hundred pints two thousand four hundred pints

OpenStudy (jdoe0001):

keep in mind that there are 2 pints to a quarter

OpenStudy (anonymous):

600 pints × (1 quart / 2 pints) = 300 quarts frequency. To be similar, all the sides have to be multiplied by the same amount. 11×2=22, 14×2=28, and 22×2=44. That doesn't show up in the list. 11×3=33, 14×3=42, and 22×3=66. That shows up in the list. 33,42,66